### v1

# dlordinal: a Python package for deep ordinal classification

Abstract
dlordinal is a new Python library that unifies many recent deep ordinal classification methodologies available in the literature. Developed using PyTorch as underlying framework, it implements the top performing state-of-the-art deep learning techniques for ordinal classification problems. Ordinal approaches are designed to leverage the ordering information present in the target variable. Specifically, it includes loss functions, various output layers, dropout techniques, soft labelling methodologies, and other classification strategies, all of which are appropriately designed to incorporate the ordinal information. Furthermore, as the performance metrics to assess novel proposals in ordinal classification depend on the distance between target and predicted classes in the ordinal scale, suitable ordinal evaluation metrics are also included. dlordinal is distributed under the BSD-3-Clause license and is available at https://github.com/ayrna/dlordinal.
Keywords: Deep learning, Ordinal classification, Python, PyTorch, Soft Labelling
1 Introduction
The term ordinal classification (also known as ordinal regression) refers to supervised learning problems characterised by an inherent order relationship between the class labels (Vargas et al., 2023b; Gutiérrez and García, 2016). For example, when classifying the nature of a tumour lesion into varying degrees of severity, the set of possible labels might be benign, premalignant (precancerous), in situ, malignant (cancerous). This scenario exemplifies an ordinal classification problem, as it involves distinct degrees of a condition. The primary objective of ordinal classification is to minimise the magnitude of misclassification errors while considering the inherent order between the labels of the problem.
This is accomplished by considering the varying costs associated with different types of errors. Specifically, classification errors between adjacent classes are less significant than errors between distant classes. Thus, in the former example, the error committed when misclassifying a benign tumour as malignant should be greater than the error incurred when misclassifying it as premalignant (precancerous). Therefore, ordinal classification approaches are designed to account for the severity of misclassification errors by incorporating the ordinal nature of the labels into both the learning algorithm and the evaluation metrics.
Some existing packages, such as ORCA (Sánchez-Monedero et al., 2019), in Matlab, or mord (Pedregosa-Izquierdo, 2015), in Python, implement ordinal classification approaches for machine learning. Moreover, some previous works have released their source code [^1] https://github.com/rpmcruz/unimodal-ordinal-regression, as is the case of some unimodal distributions for deep ordinal classification tasks (Cardoso et al., 2023). Nevertheless, these works do not provide a structured Python package, so they can not be easily installed and used. In this way, to the best of the authors’ knowledge, there is not a framework that bridges the gap between ordinal classification and Deep Learning (DL), providing the researchers with a single framework integrating numerous ordinal approaches and other functionalities, useful for developing novel methodologies.
In this paper, we introduce dlordinal, a Python toolkit that implements ordinal classification methodologies for DL developed on top of the open-source PyTorch framework. Our software offers a wide number of DL approaches that leverage ordinal information inherent in the problem domain. The documentation and code of the dlordinal toolkit is available at https://github.com/ayrna/dlordinal.
2 Code design and features
The dlordinal package is designed with a modular architecture, facilitating its use and extension. It perfectly integrates with third-party software such as skorch (Tietz et al., 2017) for a wide range of classification tasks. The dependency level is kept to a minimum, relying on essential libraries including scikit-learn (Pedregosa et al., 2011), numpy (Harris et al., 2020), PyTorch (Paszke et al., 2019), pandas (McKinney et al., 2010), scipy (Virtanen et al., 2020), scikit-image (Van der Walt et al., 2014), tqdm (da Costa-Luis, 2019), and pillow. Additionally, dlordinal supports downloading and loading common ordinal image benchmark datasets, such as FGNet (Fu et al., 2015) or Adience (Eidinger et al., 2014). This facilitates easy benchmarking of novel ordinal approaches.
Figure 1: Representation of the classes and functions included in the various modules of the dlordinal package.
As mentioned above, the dlordinal package features a modular design (see Figure 1), divided into several clearly defined modules:
  • Datasets: this module provides methods for loading the FGNet (Panis et al., 2016) and Adience (Eidinger et al., 2014), two ordinal datasets for benchmarking purposes. These datasets are useful for evaluating the performance of models in facial age estimation, a common ordinal classification task. FGNet offers a comprehensive collection of facial images across different ages, providing a valuable resource for studying facial ageing. Similarly, the Adience dataset includes unfiltered images collected from various sources, making it suitable for robust age and gender estimation. The tutorials included in dlordinal demonstrate how to effectively utilise these datasets for model training and evaluation.
  • Dropout: this module includes a new hybrid dropout methodology (Bérchez-Moreno et al., 2024) modifying the standard dropout approach (Srivastava et al., 2014) to consider the order information of the target. In this way, the conventional dropout technique randomly drops neurons during training to prevent overfitting and improve generalisation, while the ordinal version analyses correlation between neurons and targets to promote those neurons that contribute to a better ordering of the classes.
  • Output layers: dlordinal implements the output layers proposed in the current literature. It includes an implementation of the Cumulative Link Models (CLM) output layer (Vargas et al., 2020), which uses a projection obtained by an arbitrary deep neural network model. It also includes an implementation of the stick-breaking strategy proposed in Liu et al. (2020), which computes the class probabilities considering the ordinal information. Lastly, it implements an ordinal fully connected layer based on the Ordinal Binary Decomposition (OBD) approach described in Barbero-Gómez et al. (2023), which decompose the ordinal problem in several binary classification problems while preserving the ordinal information.
  • Soft labelling: this module implements several functions for generating soft labels using different probability distributions. This module forms the basis of the soft labelling strategy and is used in the corresponding loss functions. Soft labelling approaches replace one-hot encoded targets with the objective of regularising model predictions. The underlying distributions generating these soft labels should be unimodal when dealing with ordinal classification. Specifically, for an ordinal classification problem with a given number of categories, the module includes state-of-the-art methodologies for generating soft labels using several distributions, such as Poisson or binomial (Liu et al., 2020), exponential (Vargas et al., 2023c), beta (Vargas et al., 2022), and triangular (Vargas et al., 2023b, a).
  • Losses: several state-of-the-art ordinal loss functions are implemented in dlordinal. For example, regularised categorical cross-entropy loss functions based on the soft labelling strategy, considering any of the functions mentioned in the previous module. The weighted kappa loss function proposed in de La Torre et al. (2018) is also included. The latter is an ordinal loss function based on a continuous version of the weighted kappa metric. Moreover, for the OBD approach (Barbero-Gómez et al., 2023), a specific loss function is also implemented. This loss function considers the error of all the binary problems by computing the Mean Square Error (MSE) loss between the output of the model and the ideal output vector for each class.
  • Wrappers: this module provides wrappers that enable the adaptation of any existing methodology to ordinal classification. An example can be found for the method introduced in Barbero-Gómez et al. (2023). It integrates any type of architecture into an OBD framework, and enhances the learning process for ordinal problems utilising the Ordinal Error-Correcting Output Codes (ECOC) distance loss (Barbero-Gómez et al., 2023).
  • Metrics: in this module, several ordinal classification metrics are included: the 1-off accuracy, which measures the proportion of predictions that are either correct or off by one category. The Average Mean Absolute Error (AMAE) (Baccianella et al., 2009) and the Maximum Mean Absolute Error (MMAE) (Cruz-Ramírez et al., 2014). Both are robust measures that consider the imbalance degree present in the problem by averaging mean absolute errors or taking the maximum of the mean absolute errors computed for each class, respectively. The Quadratic Weighted Kappa (QWK), based on the weighted kappa index, aims to quantify the inter-rater agreement on classifying elements into a set of ordered categories. The QWK is specially appropriate to measure how well the model is predicting extreme classes (de La Torre et al., 2018). The Ranked Probability Score (RPS), presented in Epstein (1969), measures the discrepancy between the predicted cumulative distribution and the actual cumulative distribution of the ordinal target. Finally, the Geometric Mean of the Sensitivities of the Extreme Classes (GMSEC) measures classification performance by focusing solely on the extreme classes, which are often the most important in certain ordinal problems.
3 Sample code
An example demonstrating the usage of some of the ordinal elements of the dlordinal toolkit, including a dataset, a loss function and some metrics, is as follows:
1import numpy as np
2from dlordinal.datasets import FGNet
3from dlordinal.losses import TriangularCrossEntropyLoss
4from dlordinal.metrics import amae, mmae
5from skorch import NeuralNetClassifier
6from torch import $n$
7from torch.optim import Adam
8from torchvision import models
9from torchvision.transforms import Compose, ToTensor
10
11# Download the FGNet dataset
12fgnet_train = FGNet(
13 root="./datasets",
14 train=True,
15 target_transform=np.array,
16 transform=Compose([ToTensor()]),
17)
18fgnet_test = FGNet(
19 root="./datasets",
20 train=False,
21 target_transform=np.array,
22 transform=Compose([ToTensor()]),
23)
24
25num_classes_fgnet = len(fgnet_train.classes)
26
27# Model
28model = models.resnet18(weights="IMAGENET1K_V1")
29model.fc = $n$.Linear(model.fc.in_features, num_classes_fgnet)
30
31# Loss function
32loss_fn = TriangularCrossEntropyLoss(num_classes=num_classes_fgnet)
33
34# Skorch estimator
35estimator = NeuralNetClassifier(
36 module=model,
37 criterion=loss_fn,
38 optimizer=Adam,
39 lr=1e-3,
40 max_epochs=25,
41)
42
43estimator.fit(X=fgnet_train, y=fgnet_train.targets)
44train_probs = estimator.predict_proba(fgnet_train)
45test_probs = estimator.predict_proba(fgnet_test)
46
47# Metrics
48amae_metric = amae(np.array(fgnet_test.targets), test_probs)
49mmae_metric = mmae(np.array(fgnet_test.targets), test_probs)
50print(f"Test AMAE: {amae_metric}, Test MMAE: {mmae_metric}")
4 Conclusions
dlordinal has been developed to bring the practitioners with a Python package including deep ordinal classification methodologies proposed in recent literature. It is built on top of the well-known PyTorch framework. Up-to-the-knowledge of the authors, there are no previous software packages unifying all deep ordinal classification approaches under a single framework. Consequently, the scientific community should benefit from using dlordinal to develop and test novel methodologies and to apply existing ones to real-world problems with minimal effort.
Acknowledgments and Disclosure of Funding
The present work has been supported by the “Agencia Estatal de Investigación (España)”, Spanish Ministry of Research and Innovation (grant ref.: PID2020-115454GB-C22 / AEI / 10.13039 / 501100011033), by the European Commission, project Test and Experiment Facilities for the Agri-Food Domain, AgriFoodTEF (grant ref.: DIGITAL-2022-CLOUD-AI-02, 101100622), and by the ENIA International Chair in Agriculture, University of Córdoba (grant ref.: TSI-100921-2023-3), funded by the Secretary of State for Digitalisation and Artificial Intelligence and by the European Union - Next Generation EU. Recovery, Transformation and Resilience Plan. F. Bérchez-Moreno has been supported by “Plan Propio de Investigación Submodalidad 2.2 Contratos predoctorales” of the University of Córdoba. R. Ayllón-Gavilán has been supported by the “Instituto de Salud Carlos III” (ISCIII) and EU (grant ref.: FI23/00163). D. Guijo-Rubio has been supported by the “Agencia Estatal de Investigación (España)” MCIU/AEI/10.13039/501100011033 and EU NextGenerationEU/ PRTR (grant ref.: JDC2022-048378-I).